Kim wants to determine a 80 percent confidence interval for the true proportion of high school students in the area who attend their home basketball games. How large of a sample must she have to get a margin of error less than 0.02? [If no estimate is known for p, let p^p^ = 0.5]
Answer #1

Solution :

Given that,

= 0.5

1 - = 1-0.5=0.5

margin of error = E = 0.02

Z/2 = 1.28

sample size = n = (Z / 2 / E)2 * * (1 - )

= (1.28/0.02)2 *0.5*0.5

= 1024

sample size = n = 1024